# CLIP Won’t Learn Object-Attribute Binding from Natural Data and Here is Why

### v1

@@ Abstract
Contrastive vision-language models like CLIP are used for a large variety of applications, such as zero-shot classification or as vision encoder for multi-modal models. Despite their popularity, their representations show major limitations. For instance, CLIP models learn bag-of-words representations and, as a consequence, fail to distinguish whether an image is of “a yellow submarine and a blue bus” or “a blue submarine and a yellow bus”. Previous attempts to fix this issue added hard negatives during training or modified the architecture, but failed to resolve the problem in its entirety. We suspect that the missing insights to solve the binding problem for CLIP are hidden in the arguably most important part of learning algorithms: the data. In this work, we fill this gap by rigorously identifying the influence of data properties on CLIP’s ability to learn binding using a synthetic dataset. We find that common properties of natural data such as low attribute density, incomplete captions, and the saliency bias, a tendency of human captioners to describe the object that is “most salient” to them have a detrimental effect on binding performance. In contrast to common belief, we find that neither scaling the batch size, i.e., implicitly adding more hard negatives, nor explicitly creating hard negatives enables CLIP to learn reliable binding. Only when the data expresses our identified data properties CLIP learns almost perfect binding.
@@ 1 Introduction
Contrastive Vision-Language models like CLIP [19] are a cornerstone of computer vision. CLIP-like models are used for various applications, ranging from zero-shot image classification to text-to-image and image-to-text retrieval. The image encoders of CLIP-like models are a standard choice for the vision encoder in multi-modal language models [2, 7] and the text encoders have been used to guide image generation [17, 20, 21]. To improve CLIP or mitigate deficiencies, a whole family of variants with minor architectural differences or modifications of the loss function have been proposed [1, 8, 16, 25, 27, 28], which we will refer to as CLIP-like models.
@@
Despite the great success and broad applicability of CLIP-like models, they unfortunately fail at the fundamental task of object-attribute binding [15, 23, 26]. For instance, as shown in Fig. 1(a) CLIP can’t distinguish the “yellow submarine and blue bus” and the incorrect caption in which the color is swapped. Similarly, Yuksekgonul et al. [26] found CLIP models to represent images and text in a bag-of-words (BOW) representation, i.e., CLIP appears to extract a set of concepts from an image and represents an image by an unordered set of these concepts. Naturally, a BOW representation is, for many tasks, not sufficient, can result in unexpected results, and can even cause security risks. For example, imagine a self-driving car using a BOW representation, and thus failing to bind red to the traffic light. As a result, a pedestrian wearing a green shirt crossing the street might cause the car to accelerate.
@@
Figure 1: (a) An example of the object-attribute binding problem. CLIP may score the incorrect caption, with wrong color attribution, i.e. “blue submarine and yellow bus” higher than the correct one. (b) We identify various data properties that contribute to poor object-attribute binding. As can be seen, if 1) too few or too many attributes per object are in the caption, or 2) the number of objects in image or 3) caption mentions too few objects, 4) captioners focus more on salient objects, a bag-of-words (BOW) representation is sufficient to find the correct sample within a batch (highlighted by blue squares on diagonal of similarity matrices). Here we study how these data properties influence object-attribute binding of CLIP models and find (c) setups that lead to robust binding accuracy of CLIP models, suggesting the binding problem of CLIP is a data problem.
@@
Multiple previous works tried to resolve CLIP’s inability to learn reliable object-attribute binding. For instance, Yuksekgonul et al. [26] propose that the root cause of CLIP’s failure to learn binding is that attribute binding can be ignored when training with a simple contrastive loss, as a BOW representation is sufficient to associate each caption to the correct image. As a remedy, they propose to mine hard negatives. However, as revealed later in Hsieh et al. [12], even when training with such negatives CLIP’s binding ability increases only marginally on a more general test set. Others assume that the CLIP architecture limits its ability to learn object-attribute binding and try to fix the problem by adding an object-centric learning inspired cross-modal interaction layer [1]. But these changes are also not sufficient to solve the problem entirely.
@@
This leads to the central question: What prevents CLIP from learning binding? Is it really the poor scaling properties of contrastive losses in a large combinatorial space, as indicated by Yuksekgonul et al. [26] or architectural limitations as proposed by Assouel et al. [1]? We suspect that neither fully explains it. Previous works focused on the loss and architecture, but missed arguably the most important component of each learning algorithm: the data. Indeed, as shown in Fig. 1(b), we find various data properties that are common in web-scale image-caption datasets that promote the learning of BOW representations, as they make the learning of binding obsolete. Could properties of the data itself limit the binding of CLIP models? Could better filtering or re-captioning help? And what would ideal data look like?
@@
Unfortunately, it is nontrivial to study the influence of data properties of web-scraped data: they are generally unknown and annotating them is too expensive. Without annotation of such data properties it is impossible to run clean data interventions. Even with these annotations, it is necessary to train many CLIP models on different settings, which quickly becomes too expensive. As a remedy, we define a synthetic data generating process that gives us full control over the data properties. To stay close to the statistics of real data we annotate the data properties for a few images from CC12M [6]. Diverging from this real setup, we can study the influence of individual properties.
@@
But can we be sure that relationships found on synthetic data transfer to real data? Note that studying learning algorithms in a controlled setup reveals basic properties and relations of this learning algorithm. We do not expect that the learning algorithm changes its behavior fundamentally in larger scale training on more complex data. A remaining caveat is that the relations and properties found in our setup could be overshadowed by other factors that are not modeled in our simplified setting, limiting their practical use, but even then this work still contributes to a better understanding of CLIP models.
@@
In summary, our contributions are the following: 1) We design a clean synthetic dataset which allows us to study object-attribute binding of CLIP models in a controllable full-information setup. In particular, we analyze the influence of the ratio of multi-object to single-object images, the influence of ratio of objects described in the caption, the number of attributes per object in the caption and the influence of saliency bias of annotators on object-attribute binding of CLIP models. 2) We find that the number of objects in the image and in the caption have a significant influence on binding and that the multi-object ratio on images is more important. We find an inverse u-shaped relation between attributes in the caption and binding accuracy, i.e., too few and too many attributes in the caption can both hurt binding. Further, we find that the saliency bias of annotators might be one of the main limiting factors limiting object-attribute binding. 3) We validate that our results are not an artifact of a bad choice of the experimental setup by verifying that the trend exists when scaling up the batch size, model complexity and testing out-of-distribution. 4) Finally, we show that hard negatives alone are not sufficient to learn strong binding, and that the data affects the binding accuracy more.
@@ 2 Related Works
Yuksekgonul et al. [26] show that CLIP learns a bag-of-words (BOW) representation and hypothesize this is due to batches lacking hard negatives. Without them, the contrastive loss can be minimized with a BOW representation which CLIP learns (following a simplicity bias). Similarly, Tang et al. [23] report a phenomenon termed Concept Association Bias (CAB), where CLIP learns strong associations between highly correlated factors (e.g. “purple” and “eggplant”), leading to incorrect or unexpected zero-shot prediction, and exacerbating binding failures. In summary, multiple works have shown major limitations of CLIP’s ability to bind attributes to objects [15, 23, 26], supported by the poor performance on binding benchmarks  [11, 12, 26].
@@
Multiple works aimed to improve CLIP’s binding performance [1, 9, 26]. Yuksekgonul et al. [26] propose NegCLIP, which constructs hard-negatives by swapping attributes, object and relations in the caption, then fine-tunes CLIP with them. However, later works revealed that the success was overestimated, due to overfitting to specific types of binding used for testing [12]. Assouel et al. [1] try to resolve the problem via two interventions: 1) adding cross-modal interactions, inspired by object-centric learning, and 2) using an LLM to decompose captions into sub-captions of objects and their attributes or relations. Despite outsourcing most of the binding to an LLM their approach only partially solves the binding problem, as can be seen by the reported binding-accuracy on Sugarcrepe [12]. Koishigarina et al. [13] argue that the binding information is present unimodally and is lost during the cosine similarity computation and propose a contrastively trained transformation matrix to linearly align representations before similarity computation, showing decent but partial improvements.
Enriching the captions with LLMs also leads to some improvements [9, 10], as does generating hard negatives via in-context learning [18], however, these methods also do not solve the binding problem entirely. We suspect, that these approaches are limited by the limited understanding how properties of the training data influence binding.
In summary, previous efforts identified various factors and tricks for improving CLIP on binding benchmarks. Only a few [26] also take the data into account, usually limited to hard negatives created following simple rules, which limits the generalization of the method [12]. In contrast to prior works, we study the binding problem of CLIP in detail from a data-centric perspective. In particular, we study the influence of common data properties of web-scraped image-caption data on the CLIP binding performance, and show that these data properties have significant influence on binding and CLIP trained in a favorable data setting learns almost perfect binding.
@@ 3 Experimental Setup
To systematically study the influence of data properties on object-attribute binding we require precise control over data characteristics, which natural datasets cannot provide. As a remedy, we create a synthetic dataset that enables explicit manipulation of these data properties.
@@ 3.1 MADMAN: Fully Controllable Synthetic Dataset
Figure 2: MADMAN example and evaluation protocols. To evaluate recognition we create zero-shot classification prompts by changing the target attribute. To evaluate binding we swap the target attribute to create a negative and perform zero-shot classification with the true and negative caption.
@@
Our dataset, Multi Attribute and Digit for Multi-Attribute biNding (MADMAN), is designed to study object attribute binding in a controlled way. A dataset for this purpose needs to entail images with multiple objects, each with a set of attributes, all of which are described in a caption. We use the MAD [22] dataset as the starting point, add support for multiple objects per image, and extend it with more attributes. We show an example image with its caption in Fig. 2 and in the Appendix, section 5.1.
Following MAD, we use MNIST [14] digits as object classes and use augmentations from Castro et al. [5] to obtain attributes. We use the same augmentations as used for MAD [22] and add rotation to end up with six types of attributes in total. For multi-object images, we create a 3x3 grid, randomly filling 2 cells with images and the rest with background color. Captions are created by chaining object classes with their attributes, using “and” to separate the objects (see Appendix, section 5.1 for a complete list of augmentations and examples from our dataset).
[…]
3.2 Evaluation Metrics
@@
Recognition. Our final goal is to evaluate object-attribute binding, however, a prerequisite for binding is that the attribute in question can be recognized.
We evaluate recognition accuracy per object and attribute. For each object $o$ and attribute $a$, we create ground truth candidate prompts and vary only $a$. We provide an overview in Fig. 2.
@@
    Filter out attributes $a$ for which recognition is below $p_{\mathrm{chance}}$
@@
    Evaluate swap-accuracy for the remaining samples
Criterion 1) is necessary, as the attribute binding metric has no meaning if the attribute is not even recognized. To filter out cases where performance is close to chance, we filter based on $1.1\times p_{\mathrm{chance}}$ for each attribute. Criterion 2) factors out failures of binding that are caused by a failed recognition. Finally, for the remaining samples we compute swap-accuracy [12, 26]:
[…]
where $c_{n}^{\text{gt}}$ is the ground truth caption $c$ of sample $n$ and $c_{n}^{\text{swap}}$ is derived from $c_{n}^{\text{gt}}$ by swapping attribute $a$ between the objects and $p(c)$ indicates the probability the model assigns to caption $c$ belonging to the image. As this is a two-alternative forced choice task, the chance level is 0.5. We train each model with three random seeds and report the mean. We plot the mean and the 95% confidence interval.
Determining realistic parameters for the data properties. To build a synthetic dataset with roughly similar statistics as real data we annotate the 4 data properties studied in this work for 100 images from CC12M [6]. For details on the annotation refer to the appendix 5.2.
Defining Out-of-Distribution. To evaluate generalization capabilities, we define the following combinations as out-of-distribution (OOD):
  • color: $\{green,red\}\times\{0,3\}$, $\{blue,magenta\}\times\{4,5\}$
  • scaling: $\{large\}\times\{3,7\}$, $\{small\}\times\{4,9\}$
and include them only in our OOD testset.
Experimental setup for the model. Compared to natural datasets, MADMAN exhibits only few factors of variation. Both the number of object classes and attributes are limited. Accordingly, we adjust our model and batch size to match this smaller scale. Batch size plays a critical role in contrastive learning models, as it directly affects the likelihood of sampling hard negatives within a batch. This, in turn, influences the number of distinct features (like binding) the model must learn to minimize the loss. To balance this out we reduce the batch size accordingly: We use a batch size of 16 and an embedding size of 32. We validate these choices in section 4.3, demonstrating that the observed trends are robust to variations in the specific configuration. Both the text and image encoders are small transformers, each with 6 layers and 4 attention heads.

### v2

# Common Data Properties Limit Object-Attribute Binding in CLIP

@@ Abstract
Contrastive vision-language models like CLIP are used for a large variety of applications, such as zero-shot classification or as vision encoder for multi-modal models. Despite their popularity, their representations show major limitations. For instance, CLIP models learn bag-of-words representations and, as a consequence, fail to distinguish whether an image is of “a yellow submarine and a blue bus” or “a blue submarine and a yellow bus”. Previous attempts to fix this issue added hard negatives during training or modified the architecture, but failed to resolve the problem in its entirety. We suspect that the missing insights to solve the binding problem for CLIP are hidden in arguably the most important part of learning algorithms: the data. In this work, we fill this gap by rigorously identifying the influence of data properties on CLIP’s ability to learn binding using a synthetic dataset. We find that common properties of natural data such as low attribute density, incomplete captions, and the saliency bias, a tendency of human captioners to describe the object that is “most salient” to them, have a detrimental effect on binding performance. In contrast to common belief, we find that neither scaling the batch size, i.e., implicitly adding more hard negatives, nor explicitly creating hard negatives enables CLIP to learn reliable binding. Only when the data expresses our identified data properties does CLIP learn almost perfect binding.
@@ 1 Introduction
Contrastive Vision-Language models like CLIP [22] are a cornerstone of computer vision. CLIP-like models are used for various applications, ranging from zero-shot classification to text-to-image and image-to-text retrieval. The image encoders of CLIP-like models are a standard choice for the vision encoder in multi-modal language models [2, 8] and the text encoders have been used to guide image generation [20, 23, 24]. To mitigate deficiencies of CLIP, a whole family of variants with minor architectural differences or modifications of the loss function have been proposed [1, 9, 18, 28, 30, 31], which we will refer to as CLIP-like models.
@@
Despite the great success and broad applicability of CLIP-like models, they unfortunately fail at the fundamental task of object-attribute binding [17, 26, 29]. For instance, as shown in Fig.˜1(a) CLIP can’t distinguish the “yellow submarine and blue bus" and the incorrect caption in which the color is swapped. This is linked to the finding by Yuksekgonul et al. [29] that CLIP models represent images and text in a bag-of-words (BOW) representation, i.e., CLIP appears to extract a set of concepts from an image and represents an image by an unordered set of these concepts. Naturally, a BOW representation is, for many tasks, insufficient, can result in unexpected results, and can even cause security risks. For example, a self-driving car using BOW representation might fail to bind red to the traffic light and confuse a green-shirted pedestrian with a green traffic light.
@@
Figure 1: (a) The object-attribute binding problem: CLIP can’t distinguish between two captions with swapped attributes. (b) We identify data properties that contribute to poor object-attribute binding. If 1) too few or too many attributes per object are in the caption, or 2) too few images with multiple objects or 3) caption mentions too few objects, 4) captioners focus more on salient objects, a bag-of-words (BOW) representation is sufficient to find the correct sample within a batch. Here we study how these data properties influence object-attribute binding of CLIP models and find (c) setups that lead to robust binding accuracy of CLIP models, suggesting the binding problem of CLIP is a data problem.
@@
Multiple previous works tried to resolve CLIP’s inability to learn reliable object-attribute binding. For instance, Yuksekgonul et al. [29] claim that CLIP fails to learn binding because a BOW representation is sufficient to associate each caption to the correct image. As a remedy, they propose to mine hard negatives. However, as revealed later in Hsieh et al. [13], CLIP’s binding ability increases only marginally on a more general test set. Others assume that CLIP’s architecture limits its ability to learn object-attribute binding and try to fix the problem by adding an object-centric learning inspired cross-modal interaction layer [1]. But these changes are also insufficient to solve the problem entirely.
@@
This leads to the central question: What prevents CLIP from learning binding? Is it the poor scaling properties of contrastive losses in a large combinatorial space, as indicated by Yuksekgonul et al. [29], or architectural limitations as proposed by Assouel et al. [1]? We suspect that neither fully explains it. Previous works focused on the loss and architecture, but missed arguably the most important component of each learning algorithm: the data. Indeed, as shown in Fig.˜1(b), we find various data properties common in image-caption datasets to promote BOW representations, as they make the learning of binding obsolete. Could properties of the data itself limit the binding of CLIP models? Could better filtering or re-captioning help? And what would ideal data look like?
@@
Unfortunately, studying the influence of data properties is nontrivial for web-scraped data: the properties are unknown and annotating them is too expensive. Without this annotation it is impossible to run clean data interventions. Even with these annotations, it is necessary to train many CLIP models on different settings, which quickly becomes too expensive. As a remedy, we define a synthetic data generating process that gives us full control over the data properties. To ensure the data properties of our synthetic dataset are representative of real data we annotate the data properties for 100 images from CC12M [7] and pick our default setup close to those observed on real data. Diverging from this “realistic” setup, we can study the influence of individual properties.
@@
But can we be sure that relationships found on synthetic data transfer to real data? Note that studying learning algorithms in a controlled setup reveals basic properties and relations of this learning algorithm. We do not expect that the learning algorithm changes its behavior fundamentally in larger scale training on more complex data. A remaining caveat is that the relations found in our setup could be overshadowed by other factors on real data. Even then this work still contributes to a better understanding of CLIP models.
@@
In summary, our contributions are the following: 1) We design a clean synthetic dataset which allows us to study object-attribute binding of CLIP models in a controllable full-information setup. In particular, we analyze the influence of the ratio of multi-object to single-object images, the influence of number of objects described in the caption, the number of attributes per object in the caption, and the influence of saliency bias of annotators on object-attribute binding of CLIP models. 2) We find that the number of objects in the image and in the caption have a significant influence on binding. We find an inverse u-shaped relation between attributes in the caption and binding accuracy, i.e., too few and too many attributes in the caption both hurt binding. We further find that the saliency bias of annotators might be a key factor limiting object-attribute binding. 3) We confirm that our findings are not due to flaws in the experimental design by scaling up the batch size and model complexity. Furthermore, we observe that the trends persist out-of-distribution, ruling out that our models learned a shortcut. 4) Finally, we show that hard negatives alone are insufficient to learn strong binding, and that the data affects the binding accuracy more.
@@ 2 Related Works
Yuksekgonul et al. [29] show that CLIP learns a bag-of-words (BOW) representation and hypothesize this is due to batches lacking hard negatives. Without them, the contrastive loss can be minimized with a BOW representation which CLIP learns (following a simplicity bias). Similarly, Tang et al. [26] identify Concept Association Bias (CAB), where CLIP learns strong associations between highly correlated factors (e.g. “purple” and “eggplant”) also leading to binding failures.
@@
Several approaches aimed to improve CLIP’s binding performance [1, 10, 29]. Yuksekgonul et al. [29] propose NegCLIP, which is fine-tuned on hard negatives constructed by swapping attributes, object, and relations in the caption. However, later works revealed that the success was overestimated, due to overfitting to specific types of binding used for testing [13]. Assouel et al. [1] propose cross-modal interactions, inspired by object-centric learning, and LLM-based caption decomposition. Despite outsourcing most of the binding to an LLM their approach only partially solves the binding problem, as can be seen by the reported binding-accuracy on Sugarcrepe [13]. Enriching the captions with LLMs also leads to some improvements [10, 11], as does generating hard negatives via in-context learning [21], however, these methods also do not solve the binding problem entirely. Koishigarina et al. [15] argue that the binding information is present unimodally and is lost during the cosine similarity computation and propose a contrastively-trained transformation matrix for alignment, showing partial improvements. However, their study centers on overly simplistic datasets.
The datasets used to study object-attribute binding in a controlled setting are based on CLEVR [14] in [17] and PUG:SPAR [3] in [15]. However, they are limited by their low complexity. Note, that both datasets are rendered images, however, complexity for the object-attribute binding does not arise from image quality and realism. Instead, the number of attributes and their combinations with objects is the key driver of the binding task complexity. For instance, PUG:SPAR only has one attribute type (color/material). CLEVR [14] only has 3 object types with 3 attributes, 2 of which have only 2 possible values. As such, these datasets are not sufficiently complex and do not allow for enough control over their data properties to systematically study the influence of training data on CLIP’s binding performance. For evaluation of models trained on large scale image-caption datasets benchmarks from the Crepe family [12, 13, 19] are commonly used. Unfortunately, they do not provide the required control for a systematic analysis of the influence of data properties on binding performance.
In contrast to prior work, we take a data-centric approach to study CLIP’s binding problem in detail and develop a synthetic dataset that enables systematic investigation of how common properties of web-scraped image-caption data influence CLIP binding performance. We demonstrate that these data properties of the training data significantly impact binding, and that CLIP trained under favorable data conditions achieves high binding performance.
@@ 3 Experimental Setup
To systematically study the influence of training data properties on object-attribute binding we require precise control over data characteristics, which natural datasets cannot provide. As a remedy, we create a synthetic dataset that enables explicit manipulation of these data properties.
@@ 3.1 MADMAN: Fully Controllable Synthetic Dataset
Figure 2: MADMAN example and evaluation protocols. To evaluate recognition, we create zero-shot classification prompts by changing the target attribute. To evaluate binding, we swap the target attribute to create a negative and perform zero-shot classification with the true and negative caption. In practice, we evaluate on captions with 3 and 4 attributes per object.
@@
We show an example of our dataset Multi-Attribute and Digit for Multi-Attribute biNding (MADMAN) in Fig.˜2 and in the Appendix, Section˜5.1. It is designed to study object-attribute binding in a controlled way. A dataset for this purpose needs to entail images with multiple objects, each with a set of attributes, all of which are described in a caption. We use the MAD [25] dataset as the starting point, add support for multiple objects per image, and extend it with rotation. For multi-object images, we create a 3x3 black grid, and fill 2 random cells with images. Captions are created by chaining object classes with their attributes, using “and” to separate the objects (see Appendix, Section˜5.1 for a complete list of augmentations and examples from our dataset).
3.1.1 Data properties studied with MADMAN.
To study the influence of data properties on binding, we create different versions of MADMAN controlling: 1) Two-object-in-image-probability $p(\text{two-obj-img})$, which controls the probability of a sample to contain 2 objects, 2) Two-object-in-caption-probability $p(\text{two-obj-cap}|\text{two-obj-img}=\text{True})$, which controls the probability that both objects are described in the caption. Note that this is a conditional probability (given the image is a two-object image) and not the prior probability. However, for most experiments it is identical, as we use $p(\text{multi-obj-img})=1$ in our base setup. We also study 3) the influence of the number of Attributes-per-object-in-caption. We observe that humans mention only a few attributes per object in a caption (see Appendix, Section˜5.2). To model this, we change the probability of an attribute being included in the caption. To avoid overfitting to a specific caption length we define probability distributions over the number of attributes per object, and use the expected value to distinguish them (see Appendix, Fig.˜7 to see the underlying distributions). Finally, we are interested in the effect of 4) Saliency bias. Saliency describes the properties of “parts popping out” to human attention. Images commonly contain a number of salient objects, which are more likely to be described in the caption. We model it via a center bias, i.e., we define an object to be salient if it is in the center of the image. Furthermore, salient objects in our setup are always mentioned in the caption and are mentioned first. We control $p(\text{saliency})$ which refers to the probability of an image to contain a salient object.
Determining realistic parameters for the data properties. To draw conclusions about natural data from synthetic data, we need a MADMAN version that has similar data properties as natural data. To this end, we annotate the four data properties defined above for 100 images from CC12M [7]. For details on the annotation, Appendix 5.2. We refer to this MADMAN variant as “Realistic”.
Defining Out-of-Distribution. To evaluate whether our model learns a compositional representation we need an out-of-distribution (OOD) test set. We define the following combinations as OOD:
  • color: $\{green,red\}\times\{0,3\}$, $\{blue,magenta\}\times\{4,5\}$
  • scaling: $\{large\}\times\{3,7\}$, $\{small\}\times\{4,9\}$
Here, $\times$ denotes the Cartesian product. These combinations are included only in our OOD testset (only Section˜4.4). The remaining combinations are used for training and our standard test-set.
3.2 Evaluation and Training Setup
The evaluation images always show two objects and both are described in the caption. The evaluation captions have 3 attributes per object in 50% of samples and 4 for the remaining samples. This ensures a fair comparison across different Attributes-per-object-in-caption settings, as all of them have the same probability mass for 3 and 4 attributes (see Fig.˜7). We provide an overview of our recognition and binding evaluation in Fig.˜2.
@@ 3.2 Evaluation and Training Setup
Recognition. Our final goal is to evaluate object-attribute binding, however, a prerequisite for binding is that the attribute in question can be recognized. We evaluate recognition accuracy per object and attribute. For each object $o$ and attribute $a$, we use the evaluation captions and vary only $a$ (or $o$) over all possible values to create false captions. Recognition is evaluated as zero-shot accuracy over the set of these false captions and the ground-truth caption.
@@
    Filter out attributes $a$ for which recognition is close to or below $p_{\mathrm{chance}}$
@@
    Evaluate binding-accuracy for the remaining samples
Criterion 1) is necessary, as the attribute binding metric has no meaning if the attribute is not even recognized. To filter out cases where performance is slightly above chance, we use a slightly stricter threshold: $1.1\times p_{\mathrm{chance}}$. Criterion 2) factors out failures of binding that are caused by a failed recognition. Finally, for the remaining samples we compute binding-accuracy [13, 29]:
$\text{binding-accuracy}(a)=\frac{1}{N}\sum_{n=1}^{N}p(c_{n}^{\text{gt}})>p(c_{n}^{\text{swap}_{i}}),$ (1)
where $c_{n}^{\text{gt}}$ is the ground truth caption $c$ of sample $n$, $c_{n}^{\text{swap}}$ is derived from $c_{n}^{\text{gt}}$ by swapping attribute $a$ between the objects and $p(c)$ indicates the probability the model assigns to caption $c$ belonging to the image. As this is a two-alternative forced choice task, the chance level is 0.5.
3.2.1 Model training setup.
For all results reported in Section˜4, we train small CLIP models from scratch. We train each model with three random seeds and report/plot the mean and the 95% confidence interval. We vary the data properties of MADMAN as detailed for each experiment.
Compared to natural datasets, MADMAN exhibits only a few factors of variation. Both the number of object classes and attributes are limited. Accordingly, we adjust our model and batch size to match this smaller scale. Text and image encoders are small transformers with 6 layers and 4 attention heads, each. We use a batch size of 16 and an embedding size of 32. In particular the choice of the batch size plays a critical role in contrastive learning models, as it directly affects the likelihood of sampling hard negatives within a batch. This, in turn, influences the number of distinct features (like binding) the model must learn to minimize the loss. We validate our choices in Section˜4.3, demonstrating that the observed trends are robust to variations in the specific configuration.